Tami Tyler opened Tami’s Creations, Inc., a small manufacturing company, at the beginning of the year. Getting the company through its first quarter of operations placed a considerable strain on Ms. Tyler’s personal finances. The following income statement for the first quarter was prepared by a friend who has just completed a course in managerial accounting at State University. Tami’s Creations, Inc. Income Statement For the Quarter Ended March 31 Sales (28,500 units) $ 1,140,000 Variable expenses: Variable cost of goods sold $ 461,700 Variable selling and administrative 193,800 655,500 Contribution margin 484,500 Fixed expenses: Fixed manufacturing overhead 283,500 Fixed selling and administrative 214,500 498,000 Net operating loss $ ( 13,500) Ms. Tyler is discouraged over the loss shown for the quarter, particularly because she had planned to use the statement as support for a bank loan. Another friend, a CPA, insists that the company should be using absorption costing rather than variable costing and argues that if absorption costing had been used the company probably would have reported at least some profit for the quarter. At this point, Ms. Tyler is manufacturing only one product—a swimsuit. Production and cost data relating to the swimsuit for the first quarter follow: Units produced 31,500 Units sold 28,500 Variable costs per unit: Direct materials $ 7.20 Direct labor $ 7.40 Variable manufacturing overhead $ 1.60 Variable selling and administrative $ 6.80 Required: 1. Complete the following: a. Compute the unit product cost under absorption costing. b. What is the company’s absorption costing net operating income (loss) for the quarter? c. Reconcile the variable and absorption costing net operating income (loss) figures. 3. During the second quarter of operations, the company again produced 31,500 units but sold 34,500 units. (Assume no change in total fixed costs.) a. What is the company’s variable costing net operating income (loss) for the second quarter? b. What is the company’s absorption costing net operating income (loss) for the second quarter? c. Reconcile the variable costing and absorption costing net operating incomes for the second quarter. Garrison 16e Rechecks 2017-05-04, 2017-12-06, 2018-08-21, 06_18_2019_QC_CS-165659 rev: 03_26_2019_QC_CS-163919, 10_22_2019_QC_CS-187009
Answer #1

1 a). computing unit product cost under absorption costing:

Formula for Absorption costing as follows:

Absorption cost = Direct material cost+ Direct labour cost+ Variable manufacturing overhead+Fixed manufacturing overhead

formula for Fixed manufacturing overhead per unit = Fixed manufacturing overhead cost divided by unit produced

i.e as follows Fixed manufacturing overhead cost in 1st Quater is $ 283,500

Unit produced in 1st Quater is 31500 units

Fixed manufacturing cost per unit = 283500/31500 = $ 9

Absorption cost per unit = Direct material cost+ Direct labour cost+ Variable manufacturing overhead+Fixed manufacturing overhead

= $ 7.2+$ 7.4+$ 1.6+$ 9

= $ 25.2

product cost under absorption costing is $ 25.2

1.b) The company’s absorption costing net operating income (loss) for the quarter

Net income( loss) formulae under absorption costing :

Net income = (Sales subract cost of goods sold) results to Gross margin from gross margin we have to subract selling and adminstrative expenses

As per given details newly start up company asumming opening inventories balance is nil and closing inventory balance is units produces less units sold i.e 31500 - 28500 = 3000 units

Manufacturing cost for 1st quater = Direct material cost + Direct labour cost + variable manufacturing overheads + Fixed manufacturing overheads

i.e = ( $ 7.2 * 31500) + (  $ 7.4 * 31500) ( $ 1.6 * 31500)+ $ 283500

=  $ 226800+ $ 233100+ $ 50400 +  $ 283500

Manufacturing cost for 1st quater    =  $ 793800

Cost of good sold = Units sold * Product cost

= 28500 * $ 25.2 = $ 718200

  

Net income ( Loss) under absorption costing method:

Gross profit = Sales less Cost of goods sold = $ 1,140,000 - $ 718200 = $ 421800

Net income ( Loss) = Gross profit less variable & fixed selling & Adminstration expenses = $ 421800 - $ 193800 - $ 214500

Net income = $ 13500

1.c) Reconcile the variable and absorption costing net operating income (loss):

Net income under absorption costing can be reconciled with net income under variable costing by (a) subtracting the manufacturing overheads carried forward (absorbed by closing inventories) and (b) adding the manufacturing overheads brought in (absorbed by opening inventories).

Net income (absorption costing)   

less :Fixed manufacturing overheads carried forward (closing inventories)

Add: Fixed manufacturing overheads brought in (opening inventories)

equals : Net income (variable costing)

Fixed manufacturing overheads included in closing inventories = Fixed manufacturing cost divided by units produced multiply with closing stock units

Fixed manufacturing overheads included in closing inventories = $283500/31,500 × 3,000 = $ 27000

Net income (absorption costing) = $ 13500

less :Fixed manufacturing overheads carried forward (closing inventories) = $ 27000

Add: Fixed manufacturing overheads brought in (opening inventories) = 0

equals : Net Loss (variable costing) = $ 13500

3. a) The company’s variable costing net operating income (loss) for the second quarter

Variable production costs per unit =

Direct material cost+ Direct labour cost+ Variable manufacturing overhead

= $ 7.2+$ 7.4+$ 1.6

= $ 16.2

Variable costing income statement has the following line items:

Sales = 34500* 40 = $ 1380000

less Variable production costs ( Opening goods)= 16.2 * 3000 = $ 48600

less Variable production costs = 16.2 * 31500 = $ 510300

less Variable selling & admin costs = 6.8*34500 = $ 234600

Equals Contribution margin =  $ 586500

Less Fixed manufacturing overheads = $ 283500

Less  Fixed selling & admin costs = $ 214500

equals Net income = $ 88500

Net income under variable costing =  $ 88500

3.b).The company’s absorption costing net operating income (loss) for the second quarter:

Net income ( Loss) under absorption costing method:

= Sales less Cost of goods sold less variable & fixed selling & Adminstration expenses

Sales cost per unit assuming same as in first quarter =  $ 1,140,000 / 28500 = $ 40

Sales in second quarter = 34500*$ 40 = $ 1380000

Fixed manufacturing cost per unit = $ 283500/31500 = $ 9

Absorption cost per unit = Direct material cost+ Direct labour cost+ Variable manufacturing overhead+Fixed manufacturing overhead

= $ 7.2+$ 7.4+$ 1.6+$ 9

= $ 25.2

Cost of goods sold in second quarter = Units sold * Product cost = 34500* $ 25.2 =$ 869400

Gross profit = Sales less Cost of goods sold = $ 1380000 - $ 869400 = $ 510600

Net income under absorption costing = Gross profit less variable & fixed selling & Adminstration expenses

= $ 510600 - 6.8*34500 - $ 214500

= $ 510600 - $ 234600 - $ 214500

Net income = $ 61500

3.(c) Reconcile the variable costing and absorption costing net operating incomes for the second quarter

Net income (absorption costing)   

less :Fixed manufacturing overheads carried forward (closing inventories)

Add: Fixed manufacturing overheads brought in (opening inventories)

equals : Net income (variable costing)

Opening units in second quarter = 3000

units produced in frist quarter minus units sold in first quarter = 31500 -28500 = 3000

Fixed manufacturing overheads included in closing inventories = $283500/31,500 × 0 = $ 0

Net income (absorption costing)    = $ 61500

less :Fixed manufacturing overheads carried forward (closing inventories) = $ 0

Add: Fixed manufacturing overheads brought in (opening inventories) = $283500/31,500 × 3000 = $ 27000

equals : Net income (variable costing) = $ 88500